# Visual Riddles: a Commonsense and World Knowledge Challenge for Large Vision and Language Models

### v1

@@ 1 Introduction
Humans intuitively utilize commonsense reasoning to interpret complex elements in visual scenes, a skill that current vision and language models frequently lack. For instance, the simple act of a person scratching their arm gains added significance when a mosquito is present on a nearby nightstand, as depicted in Fig. 1. While humans easily recognize such contextual nuances, existing image-understanding models struggle to integrate visual cues with world knowledge stemming from cultural aspects, life-experiences, and physical or social knowledge [1, 2, 3, 4]. This gap has spurred the development of various benchmarks like VisIT-Bench [5] and Encyclopedic VQA [6], which rely on pre-existing images to formulate challenging questions. While effective, this approach risks using images seen during pre-training of large vision-language models and restricts the scope of scenarios to those already captured, potentially limiting creativity and challenge variety.
To address these shortcoming, we introduce Visual Riddles, a benchmark comprising 400 visual riddles, each featuring a question and a synthetic image generated specifically for the challenge. The process of creating a riddle involves designing a scenario with embedded clues that appear as natural elements, then translating this scenario into both an image and a question (§2). For instance, in the top example of Fig. 1, a seemingly inconspicuous mosquito acts as a crucial clue for the source of the person’s discomfort. This method, going from scenario to image, and not the other way around, not only enhances creative liberty but also broadens the range of everyday situations explored, challenging both humans and models in their interpretative and commonsense reasoning abilities. The benchmark also incorporates textual hints and attributions (see  Fig. 1) to direct focus to clues and provide external sources of verification for answers, thus expanding the challenge’s scope and depth.
@@
Yet, as in every QA benchmark, evaluation is a key challenge in this setting. To facilitate scalable research, we introduce two additional tasks. The first is a multiple-choice version of the main task, also for the hint- and attribution-assisted variants, allowing for easy automatic scoring based on accuracy. The second task assesses the ability of models to determine the accuracy of open-ended responses in two settings: reference-free, where models evaluate responses based only on the image and the question, and reference-based, where the correct answer is also provided. This task suggests auto-raters to evaluate our riddles, and aim to advance research on such automatic evaluation methods.
Experimental results (§4) reveal a significant gap in performance between humans and state-of-the-art vision language models, with the top-performing model, Gemini-Pro-1.5 [7], only achieving 40% accuracy versus humans’ 82%. Surprisingly, the multiple-choice format proved almost as difficult, with performances slightly superior to those in the open-ended task, but with notably high performances when auxiliary data is included. Automated tests, both reference-free and reference-based, show that Gemini-Pro-1.5, the top auto-rater, matches human judgment 87% of the time in reference-based evaluations, propose a suitable method for evaluating open-ended answers. We also find that some models fail to respond to visual clues accurately when tasked with evaluating answers’ validity (§5.1). Finally, attempts to reproduce the benchmark’s images, rich in nuanced visual clues, using the same prompts and various text-to-image models (§5.2), result in a mere 15% success rate, showcasing the unique challenges visual riddles present to current generative models.
@@
Figure 2: Overview of the Visual Riddles tasks: (1) Main Task: Solve open-ended questions. (2) Utilizing Hints: Use textual aids to identify key visual clues in riddles. (3) Employing Attributions: Apply web-sourced attributions to improve world-knowledge. (4) Multiple Choice: Select the correct answer to the riddle from five options. (5) Automatic Evaluation: Evaluate open-ended answers in two scenarios— Reference-Free, assessing the correctness of a response based only on the visual riddle, and Reference-Based, comparing responses to the ground truth answer.
@@ 2 Data Collection
The Visual Riddles Challenge uses visual clues in images to test common-sense reasoning in vision-and-language models. The goal is to develop a dataset with images that represent ambiguous and culturally rich scenarios. To solve the riddles, models must detect subtle visual clues and engage in reasoning that integrates commonsense and world knowledge.
The Visual Riddles benchmark was hand-curated by seven designers experienced in generating images with text-to-image models. The design process was guided by comprehensive instructions to create visual riddles that were complex enough to challenge models yet solvable by humans. Each riddle consisted of a synthetic image paired with a question and a corresponding ground truth answer (see examples in Fig. 1). Further elaboration on the image generation process was in A.2. To generate high-quality images, the designers had access to advanced text-to-image models, including Midjourney, Ideogram, Canva, Gemini [8], SDXL [9], DALL-E 3 [10], and Stable-Diffusion [11]. The generated images included subtle clues crucial for solving the riddles, and designers were encouraged to embed cultural nuances and their own world knowledge into the riddles. The images were not only photo-realistic but also contained all necessary elements to lead to the correct answer. Designers provided additional hints to the visual clues that guided where to look in the image when trying to solve the riddle (see top example in Fig. 1) and, for riddles requiring world knowledge, included attributions to relevant sources (see bottom example in Fig. 1). After creation, each riddle was peer-reviewed by at least three other designers to ensure the hint’s clarity and the riddle’s solvability. Upon approval, the designer drafted a detailed answer that logically explained the solution based on the visual clues. The Visual Riddles benchmark is licensed under Apache License 2.0.
@@ 4.1 Models
We evaluate LLaVA-1.5-7B [12], LLaVA-1.6-34B [13], InstructBLIP-7B [14], GPT4-turbo-preview [15], Gemini-Pro-Vision [8], and Gemini-Pro-1.5 [7]. We used the most up-to-date versions of each model, as of May 2024. Each experiment required prompts of varying token lengths; notably, the prompts for attribution tasks were particularly lengthy, as they also included the attributing texts. Models selection is therefore based on their capacity to accommodate the full prompt required for each task, ensuring that inputs are processed fully without truncation. For further details, see A.4.
@@ 4.2 Open-ended VQA
We utilized Amazon Mechanical Turk to collect human open-ended answers for the benchmark’s riddles. Rigorous qualification ensured only 10 highly-rated workers, previously experienced in similar tasks, annotated each of the 400 riddles (three workers per riddle). During annotation, we instructed annotators to not only answer the question but also provide justifications. Annotators were also encouraged to utilize tools like Google Lens and different search engines to research clues and ensure accurate identification of objects. For example, in Fig. 2.4, annotators which are unfamiliar with the Kakapo bird were advised to use Google Lens to identify it. For more details on the annotation process, including full guidelines and UI screenshot examples, see A.5.
@@ 4.3 Multiple-choice VQA
Results
The results in Table 2, indicate that this multiple-choice version is comparably challenging to the open-ended visual riddles, with GPT-4, Gemini-Pro-Vision and Gemini-Pro-1.5 showing the highest accuracies of 45%, 41% and 38%, respectively. Overall, model performance on this task slightly exceeds that on the open-ended task as assessed by human evaluators. Excluding instances where models selected the “cannot determine” option enhances these figures, elevating GPT-4 to 52% and Gemini-Pro-1.5 to 48%. This trend, detailed further in A.7, reveals that models often default to “cannot determine” in the absence of sufficient information. This tendency is mitigated by providing hints and attributions, making GPT-4’s accuracy significantly increase to 69% and 82%, respectively.
@@ Results
Utilizing the Optimal Auto-Rater to Evaluate Visual Riddles
Having established Gemini-Pro-1.5 as the best auto-rater, we utilize it in a reference-based setting to assess all models on the three open-ended settings detailed in §4.2, i.e., the main task, and its hint- and attribution-assisted variants.
@@
Table 4 presents the auto-rating accuracy of various models on the open-ended task, with Gemini-Pro-1.5 as the evaluator. Gemini-Pro-1.5 achieves the highest performance, scoring 53% (40% with human rating) on open-ended questions, 62% with hints, and 29% with attributions. These results suggest that models perform better when provided with hints but not with attributions. Hints improve model performance by directing attention to visual clues, but they do not notably enhance reasoning capabilities. Conversely, when attributions are provided, models struggle to filter through irrelevant details, indicating significant challenges in solving open-ended visual riddles compared to human levels, even with auxiliary information. This highlights ongoing difficulties in improving models’ visual reasoning capabilities and bridging the gap between human and machine understanding in visual interpretation tasks.
5 Analysis
@@ 5.1 Assessing Visual Aspects of Riddles: Ablation Study with Modified Images
To assess models’ preference for specific answers, we selected 72 images from our benchmark and created modified versions, altering visual cues to invalidate ground truth answers. This study explores whether models base their answers solely on text or consider visual clues. We conducted two rounds of testing: one with original images and one with modified versions. In both instances, we presented the model with the question, the original ground truth answer, and asked,“Is the answer correct?” (see Fig. 4). Post-image modification, known ground truth answers became incorrect, make the evaluation to a binary assessment. For each model, we calculated the percentage of correctness rates on each type of image and the gap. A high gap indicates that the model identified the answer as correct for the original image and incorrect for modified one. Table 5 reveals models’ preference for specific answers despite changes in critical image elements and their struggle to identify correct answers, especially with modified images. Gemini-Pro-1.5 excels with a 60% gap, while others below 53%.
@@ 5.2 The Visual Riddles Prompt Set: A Challenge for Text-to-Image Models
6 Related Work
Our research is closely linked to commonsense reasoning in multimodal models and the evaluation of factuality in both language-only and multimodal settings.
Commonsense reasoning in multimodal models.
Recent progress in vision-language models including models like BLIP2 [21], Flamingo [22], LLaVA [12], GPT4 [15], and Gemini-Pro [8]. These developments have sparked interest in commonsense reasoning [2, 23], leading to complex visual reasoning challenges within the vision-and-language domain. These include specialized tests for understanding associations and analogies [24, 25], interpreting unusual images in WHOOPS! [26], visual abductive reasoning tasks (e.g., Sherlock; [27]), multi-modal humor understanding tasks [3], and world-instructed image-editing tasks requiring commonsense reasoning (e.g., EditWorld; [28]). A related benchmark to Visual Riddles is VisIT-Bench [5], where the authors took existing images and generated challenging questions about them. However, Visual Riddles distinguishes itself by allowing annotators more creative freedom to devise challenging scenarios and generate corresponding images and questions, rather than restricting them exclusively to what appears in the images. This also ensures the images have not been previously seen during pretraining.
Factuality evaluation.
@@ 7 Conclusions and Limitations
Our analysis of the Visual Riddles Challenge shows that state-of-the-art vision-language models face difficulties in interpreting complex visual scenarios that require commonsense reasoning. With an average success rate of 40%, models significantly lag behind human performance, which stands at 82%. This performance gap underscores the ongoing challenge of bridging human and machine understanding in complex visual interpretation tasks. Visual Riddles, though smaller than other benchmarks, was meticulously annotated to ensure a rich and diverse set of commonsense challenges. While we employed a variety of generative tools to mitigate potential biases, it is crucial to acknowledge that relying on these tools might introduce inherent limitations. Although we have made efforts to exclude visual riddles that could be considered offensive, it is still possible that individuals may find certain riddles offensive, despite our best efforts to prevent this. A promising future direction is to explore model performance across multiple images generated from the same prompt, allowing for a more robust evaluation of visual reasoning capabilities.
@@ A.2 Image Generation Guidelines
The Visual Riddles benchmark was created by seven designers experienced in generating images using text-to-image models. The annotators were instructed to create images that integrate information with world knowledge and common sense to answer a given textual question. The answer must be grounded in the data presented in the image, making it impossible to respond accurately without comprehending the image and identifying the embedded clues. To generate high-quality images, our designers are given access to advanced text-to-image models, including Midjourney [^1] https://www.midjourney.com/home, Ideogram [^2] https:///www.ideogram.ai, Canva [^3] https://www.canva.com, DALL-E 3 [10], and Stable-Diffusion [11]. Each image must be sufficiently challenging such that at least one of the evaluation models (e.g., Gemini-Pro-1.5 [7], GPT4-turbo-preview [15] and LLaVA-1.6-34B [13]) fails to correctly answer the question based on the provided image. In addition, the designers provided not only the correct answer to the question but also a hint to the image that should guide where to look in the image when trying to solve the riddle (Example in Fig. 1) and, for riddles requiring world knowledge, include attributions to relevant sources (see bottom example in Fig. 1). After creation, each image-question pair undergoes a peer review by at least three other designers to ensure the hint’s clarity and the riddle’s solvability.

### v2

@@ 1 Introduction
Humans intuitively utilize commonsense reasoning to interpret complex elements in visual scenes, a skill that current vision and language models frequently lack. For instance, the simple act of a person scratching their arm gains added context when a mosquito is present on a nearby nightstand, as depicted in Fig. 1. While humans easily recognize such contextual nuances, existing image-understanding models struggle to integrate visual cues with world knowledge stemming from cultural aspects, life-experiences, and physical or social knowledge [1, 2, 3, 4]. This gap has spurred the development of various benchmarks like VisIT-Bench [5] and Encyclopedic VQA [6], which rely on pre-existing images to formulate challenging questions. While effective, this approach risks using images seen during pre-training of large vision-language models and restricts the scope of scenarios to those already captured, potentially limiting creativity and challenge variety.
To address these shortcoming, we introduce Visual Riddles, a benchmark comprising 400 visual riddles, each featuring a question and a synthetic image generated specifically for the challenge. The process of creating a riddle involves designing a scenario with embedded clues that appear as natural elements, then translating this scenario into both an image and a question (§3). For example, as shown in the top of Fig. 1, a seemingly inconspicuous mosquito becomes a key clue to explain the person’s discomfort. This method, going from scenario to image, and not the other way around, not only enhances creative liberty but also broadens the range of everyday situations explored, challenging both humans and models in their interpretative and commonsense reasoning abilities. The benchmark also incorporates textual hints and attributions (Fig. 1) to direct focus to clues and provide external sources of verification for answers, thus expanding the challenge’s scope and depth.
@@
Yet, as in every QA benchmark, evaluation is a key challenge in this setting. To facilitate scalable research, we introduce two more tasks. The first is a multiple-choice version of the main task, including for the hint- and attribution-assisted variants, allowing for easy accuracy-based automatic scoring. The second task assesses the ability of models to determine the accuracy of open-ended responses in two settings: reference-free, where models evaluate responses based solely on the image and the question, and reference-based, where the correct answer is also given. This task suggests auto-raters to evaluate our riddles, aiming to advance research on such automatic evaluation methods.
Experimental results (§5) reveal a significant gap in performance between humans and state-of-the-art vision language models, with the top-performing model, Gemini-Pro-1.5 [7], only achieving 40% accuracy versus humans’ 82%. Surprisingly, the multiple-choice format proved nearly as challenging, yielding only slightly better results than the open-ended task; however, performance showed a significant improvement when auxiliary data, such as hints and attributions, was provided. Automated tests, both reference-free and reference-based, show that Gemini-Pro-1.5, the top auto-rater, matches human judgment 87% of the time in reference-based evaluations, thereby proposing a suitable method for automatically evaluating open-ended answers. We also find that some models fail to respond to visual clues accurately when tasked with evaluating answers’ validity (§6.1). Finally, attempts to reproduce the benchmark’s images, rich in nuanced visual clues, using the same prompts and various text-to-image models (§6.2), result in a mere 15% success rate, showcasing the unique challenges visual riddles present to current generative models.
@@
Figure 2: Overview of the Visual Riddles tasks: (1) Main Task: Solve open-ended questions. (2) Utilizing Hints: Use textual aids to identify key visual clues in riddles. (3) Employing Attributions: Apply web-sourced attributions to improve world-knowledge. (4) Multiple Choice: Select the correct answer to the riddle from five options. (5) Automatic Evaluation: Evaluate open-ended answers in two scenarios— Reference-Free, assessing the correctness of a candidate answer (CA) based only on the visual riddle, and Reference-Based, comparing CAs to the ground truth answer (GTA).
2 Related Work
Our research is closely linked to commonsense reasoning in multimodal models and the evaluation of factuality in both language-only and multimodal settings, with our benchmark uniquely focusing on fine-grained image understanding that combines commonsense and world knowledge, making direct comparisons challenging.
Commonsense reasoning in multimodal models.
Recent progress in vision-language models including models like BLIP2 [8], Flamingo [9], LLaVA [10], GPT4 [11], and Gemini-Pro [12]. These developments have sparked interest in commonsense reasoning [2, 13], leading to complex visual reasoning challenges within the vision-and-language domain. These include specialized tests for understanding associations and analogies [14, 15], interpreting unusual images in WHOOPS! [16], visual abductive reasoning tasks (e.g., Sherlock; [17]), multi-modal humor understanding tasks [3], and world-instructed image-editing tasks requiring commonsense reasoning (e.g., EditWorld; [18]). Notable benchmarks include VisIT-Bench [5] where the authors took existing images and generated challenging questions about them, OK-VQA [19] where questions require world knowledge, and others [20, 21]. However, Visual Riddles distinguishes itself by allowing annotators more creative freedom to devise challenging scenarios and generate corresponding images and questions, rather than restricting them exclusively to what appears in natural images. This also ensures the images have not been previously seen during pretraining.
Factuality evaluation.
@@ 3 Data Collection
The Visual Riddles Challenge uses visual clues in images to test common-sense reasoning in vision-and-language models. The goal is to develop a dataset with images that represent ambiguous and culturally rich scenarios. To solve the riddles, models must detect subtle visual clues and engage in reasoning that integrates commonsense and world knowledge. The Visual Riddles benchmark was hand-curated by seven designers, experienced in generating images with text-to-image models. The design process was guided by comprehensive instructions to create visual riddles that were complex enough to challenge models yet solvable by humans. Each riddle consisted of a synthetic image paired with a question and a corresponding ground truth answer (see examples in Fig. 1). To generate high-quality images, the designers had access to advanced text-to-image models, including Midjourney, Ideogram, Canva, Gemini [12], SDXL [37], DALL-E 3 [38], and Stable-Diffusion [39]. The generated images included subtle clues crucial for solving the riddles, and designers were encouraged to embed cultural nuances and their own world knowledge into the riddles. The images were not only photo-realistic but also contained all necessary elements to lead to the correct answer. Designers provided additional hints to the visual clues that guided where to look in the image when trying to solve the riddle (top example in Fig. 1) and, for riddles requiring world knowledge, included attributions to relevant sources (bottom example in Fig. 1). After creation, each riddle was peer-reviewed by at least three other designers to ensure the hint’s clarity and the riddle’s solvability. Upon approval, the designer drafted a detailed answer that logically explained the solution based on the visual clues. Further elaboration on the image generation process and the instructions for riddle creation are in A.2. The Visual Riddles benchmark is licensed under Apache License 2.0.
@@ 5.1 Models
We evaluate LLaVA-1.5-7B [10], LLaVA-1.6-34B [40], InstructBLIP-7B [41], GPT4-turbo-preview [11], Gemini-Pro-Vision [12], and Gemini-Pro-1.5 [7]. We used the most up-to-date versions of each model, as of May 2024. After submission, we included four additional models: GPT4o [42], Claude 3.5 Sonnet [43], Qwen-VL-Max [44], and Molmo-7B [45], all updated to the end of October 2024. Each experiment required prompts of varying token lengths; notably, the prompts for attribution tasks were particularly lengthy, as they also included the attributing texts. Models selection is therefore based on their capacity to accommodate the full prompt required for each task, ensuring that inputs are processed fully without truncation. For further details, see A.4.
@@ 5.2 Open-ended VQA
We utilized Amazon Mechanical Turk to collect human open-ended answers for the benchmark’s riddles. Qualification test ensured ten reliable workers, annotated each of the 400 riddles (three workers per riddle). During annotation, we instructed annotators to not only answer the question but also provide justifications. Annotators were also encouraged to utilize tools like Google Lens and different search engines to research clues and ensure accurate identification of objects. For example, in Fig. 2.4, annotators which are unfamiliar with the Kakapo bird were advised to use Google Lens to identify it. For more details on the annotation process, including full guidelines and UI screenshot examples, see A.5.
@@ Results
[…]
Caption Quality and Model Gaps
[…]
Tool Usage and Model Gaps
We analyzed all data with attribution, comprising 164 riddles (41%) where external tools, such as Google Search or Lens, could potentially aid in solving. The remaining 59% did not necessitate these tools. Model performance was assessed across both categories, revealing a slight difference: models reached 21% accuracy on riddles where external tools might have been useful and 26% on those not requiring such assistance. These findings suggest that the main difficulties for models lie in reasoning and interpreting visual clues rather than depending on external knowledge.
@@ 5.3 Multiple-choice VQA
Claude 3.5 Sonnet 46 4 48 45 -
GPT4o 55 17 67 83 -
Qwen-VL-Max 35 3 37 53 -
Molmo-7B 34 1 35 42 -
Results
The results in Table 2, indicate that this multiple-choice version is comparably challenging to the open-ended visual riddles, with GPT-4, Gemini-Pro-Vision and Gemini-Pro-1.5 showing the highest accuracies of 45%, 41% and 38%, respectively. Overall, model performance on this task slightly exceeds that on the open-ended task as assessed by human evaluators. Excluding instances where models selected the “cannot determine” option enhances these figures, elevating GPT-4 to 52% and Gemini-Pro-1.5 to 48%. This trend, detailed further in A.7, reveals that models often default to “cannot determine” in the absence of sufficient information. This tendency is mitigated by providing hints and attributions, making GPT-4’s accuracy significantly increase to 69% and 82%, respectively. Post-submission results further highlight this trend, with GPT4o leading the models at 55% accuracy overall and reaching 83% with hints.
@@ Results
Utilizing the Optimal Auto-Rater to Evaluate Visual Riddles
Having established Gemini-Pro-1.5 as the best auto-rater, we utilize it in a reference-based setting to assess all models on the three open-ended settings detailed in §5.2, i.e., the main task, and its hint- and attribution-assisted variants. Table 4 presents the auto-rating accuracy of various models on the open-ended task, with Gemini-Pro-1.5 as the evaluator. Gemini-Pro-1.5 achieves the highest performance, scoring 53% (40% with human rating) on open-ended questions, 62% with hints, and 29% with attributions. These results suggest that models perform better when provided with hints but not with attributions. Hints improve model performance by directing attention to visual clues, but they do not notably enhance reasoning capabilities. Conversely, when attributions are provided, models struggle to filter through irrelevant details, indicating significant challenges in solving open-ended visual riddles compared to human levels, even with auxiliary information. This highlights ongoing difficulties in improving models’ visual reasoning capabilities and bridging the gap between human and machine understanding in visual interpretation tasks.
@@ 6.1 Assessing Visual Aspects of Riddles: Ablation Study with Modified Images
To assess models’ preference for specific answers, we selected 72 images from our benchmark and created modified versions, altering visual cues to invalidate ground truth answers. This study explores whether models base their answers solely on text or consider visual clues. We conducted two rounds of testing: one with original images and one with modified versions. In both instances, we presented the model with the question, the original ground truth answer, and asked,“Is the answer correct?” (Fig. 5). Post-image modification, known ground truth answers became incorrect, make the evaluation to a binary assessment. For each model, we calculated the percentage of correctness rates on each type of image and the gap. A high gap indicates that the model identified the answer as correct for the original image and incorrect for modified one. Table 5 reveals models’ preference for specific answers despite changes in critical image elements and their struggle to identify correct answers, especially with modified images. Gemini-Pro-1.5 excels with a 60% gap, while others below 53%.
@@ 7 Conclusions and Limitations
Our analysis of the Visual Riddles Challenge shows that state-of-the-art vision-language models face difficulties in interpreting complex visual scenarios that require commonsense reasoning. With an average success rate of 40%, models significantly lag behind human performance, which stands at 82%. This performance gap underscores the ongoing challenge of bridging human and machine understanding in complex visual interpretation tasks. Although Visual Riddles is smaller than other benchmarks, it follows a “quality over quantity” approach, seen in recent benchmarks[51, 14, 16, 5, 52, 53], which emphasize high-quality challenges. Our dataset’s size allows for meticulous hand-curation to ensure diverse commonsense challenges, focusing on evaluation over training. While we utilized a variety of generative tools to reduce potential biases, this reliance may introduce inherent limitations, and despite efforts to exclude offensive content, some individuals may still find certain riddles inappropriate. Future work could explore automated dataset generation, creating a dedicated training set, and validating Visual Riddles as a robust test set. Generating multiple images for each prompt, coupled with repeated evaluations, would also allow more robust assessments of models’ visual reasoning capabilities.
@@ A.2 Image Generation Guidelines
The Visual Riddles benchmark was created by seven designers including four women and three men, most of whom are authors of this paper, all from the same country, experienced in generating images using text-to-image models. The annotators were instructed to create images that integrate information with world knowledge and common sense to answer a given textual question. The answer must be grounded in the data presented in the image, making it impossible to respond accurately without comprehending the image and identifying the embedded clues. To generate high-quality images, our designers are given access to advanced text-to-image models, including Midjourney [^1] https://www.midjourney.com/home, Ideogram [^2] https:///www.ideogram.ai, Canva [^3] https://www.canva.com, DALL-E 3 [38], and Stable-Diffusion [39]. Each image must be sufficiently challenging such that at least one of the evaluation models (e.g., Gemini-Pro-1.5 [7], GPT4-turbo-preview [11] and LLaVA-1.6-34B [40]) fails to correctly answer the question based on the provided image. In addition, the designers provided not only the correct answer to the question but also a hint to the image that should guide where to look in the image when trying to solve the riddle (Example in Fig. 1) and, for riddles requiring world knowledge, include attributions to relevant sources (see the full instructions in Fig. 7Fig. 8 and the bottom example in Fig. 1). After creation, each image-question pair undergoes a peer review by at least three other designers to ensure the hint’s clarity and the riddle’s solvability.
Figure 7: Guidelines for human designers to create a visual riddle — Part 1. This section includes instructions on the visual riddle creation process including requirements for captions, generative models, hints, and attribution.
Figure 8: Guidelines for human designers to create a visual riddle — Part 2. This section includes examples, and, similarly to the presented examples, designers were asked to provide additional details such as the image caption, the generative model used, a hint, and attribution for their riddles.